• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Ramesh

    Log entry time 23:49:00 on December16,2006

    Entry number 190854

    keyword=Shift Summary


    Swing Shift Summary
    (Saturday, 16th Dec, 06)






    Shift Crew:  Ramesh Subedi and Evaristo Cisbani



    Kinematics: 8

    Left HRS (electron) p0 = 2.803 GeV/c    Angle = 30.648° (main scale=31,Vernier=-30)

    Right HRS (proton) p0  = 0.832 GeV/c    Angle = 59.181° (main scale=59,Vernier=-2.8)

    Beam into the hall 83uA (MCC BCM), hall A BCM shows 89uA.



    Shift Summary






    • 16:00 Shift started with the run no 1523 passed on from the day shift




    • 16:36   Saw left hrs magnet readback frozen. Rebooted iocha14 through the web, came back to normal. This readback had been freezing somewhere between the runs 1522 and 1523 (figured out after looking at the "start of run" halog entries).




    • 19:19   Saw left hrs magnet readback frozen again. Rebooted iocha14 through the web, came back to normal.




    • 20:51   MCC took away beam for PZT test. Stopped the run 1525.



    • 21:05   PZT test finished. Beam back.







    Run Summary




    All runs are kinematics 8 production runs on 3He at 83µA with the rastered beam.


  • From the last shift Run 1523 (3.10 M)


  • 17:09 - 19:09        Run 1524 (3.08M)


  • 19:10 - 20:51        Run 1525 (2.59M)



  • 21:06 - 23:11        Run 1526 (3.00M)



  • 24:00                     Run 1527 - Passed on to the owl shift